OpenStudy (anonymous):

Solve: x+3=sqrtx^2+15 x=?

OpenStudy (anonymous):

sqrt(x^2+15) or sqrt x^2

OpenStudy (anonymous):

is important to know

OpenStudy (anonymous):

Just square both sides

OpenStudy (anonymous):

sqrt(x^2+15)

OpenStudy (anonymous):

ok than square both sides to remove the square root

OpenStudy (anonymous):

(x+3)^2 = x^2 + 15 x^2 + 6x + 9 = x^2 + 15 6x = -9 + 15 6x = 6 x = 1

OpenStudy (anonymous):

thanks